OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164

OpenStudy (amistre64):

spose you normalize it to: b0 = a19 and b2= -164

OpenStudy (amistre64):

or a21 = a19 + 2d , solve for d to find the common difference then using either value, subtract how many differences it takes to get to the first term

OpenStudy (anonymous):

Sorry my internet went out. @amistre64 So far I've gotten -53. I don't know what to do next.

OpenStudy (anonymous):

Answer choices: an = 896 - 53(n - 2) an = 896 - 53(n - 1) an = 896 + 53(n + 1) an = 896 - 53(n + 1)

OpenStudy (amistre64):

when given options, the simplest thing to do is to test them out ... which one gives us the correct values when n=19 and 21?

OpenStudy (amistre64):

they all start with 896, and all have a 53 we know we get lower so -53 is required and the proper format is that for a given n value, we assess it at the one before it, the n-1th value

OpenStudy (anonymous):

so it would be an = 896 - 53(n - 1)

OpenStudy (amistre64):

that would be the one

OpenStudy (anonymous):

896-53(20)=-164

OpenStudy (anonymous):

Thanks